(D3 - 1)y = 4 sin x.
1
Expert's answer
2020-12-20T18:03:44-0500

Given differential equation "(D^3-1)y=sinx"

Auxiliary equation corresponding to the given homogeneous differential equation is

"m^3-1=0\Rightarrow (m-1)(m^2+m+1)=0"

"\Rightarrow m=1,m=\frac{-1\pm \sqrt{1-4}}{2}=\frac{-1\pm \sqrt{3}i}{2}"

Solution to the homogeneous differential equation is

"y_{h}(x)=c_{1}e^{x}+e^{\frac{-x}{2}}\left [ c_{2}cos(\frac{\sqrt{3}x}{2})+c_{3}sin(\frac{\sqrt{3}x}{2}) \right ]"

Now let us find the particular solution using the method of undetermined coefficeints.

A particular solution using the method of undetermined coefficients is

"y_{p}(x)=Asinx+Bcosx"

"\Rightarrow y'_{p}(x)=Acosx-Bsinx, y''_{p}(x)=-Asinx-Bcosx,"

"y'''_{p}(x)=-Acosx+Bsinx"

Now "y'''_{p}(x)-y_{p}(x)=sinx"

"\Rightarrow -Acosx+Bsinx-Acosx+Bsinx=sinx"

"\Rightarrow -2Acosx+2Bsinx=sinx"

"\Rightarrow A=0, B=\frac{1}{2}"

So, the particular solution is "y_{p}(x)=\frac{1}{2}cosx"

Therefore, general solution to the given differential equation is

"y(x)=y_{h}(x)+y_{p}(x)"

That is

"y(x)=c_{1}e^{x}+e^{\frac{x}{2}}\left [ c_{2}cos(\frac{\sqrt{3}x}{2})+c_{3}sin(\frac{\sqrt{3}x}{2}) \right ]" "+\frac{1}{2}cosx"